OpenStudy (1dfanatic):

6^(-2) but it says that a negative exponent means to invert the base...idk what that means

OpenStudy (anonymous):

basicly 6\6

OpenStudy (1dfanatic):

i still don't get it...@blasterjakz

OpenStudy (anonymous):

well my tech said itmwould b3 6*-6

OpenStudy (dinamix):

\[6^{-2} = \frac{ 1 }{ 6^2 } = \frac{ 1 }{ 36 }\] @1Dfanatic

OpenStudy (anonymous):

it will be 36

OpenStudy (anonymous):

-6^2